OpenStudy (anonymous):

Use long division to find the quotient below. Enter your answer as a polynomial in descending order (x^2 + 14x + 48)/ (x + 8)

OpenStudy (anonymous):

See the links I provided in your other question. Learn the method and you will be able to do all this on your own.

OpenStudy (anonymous):

Is it x+6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this can be done easily by noting that x^2+14x+48 = x^2 + 6x+8x+48 = x(x+6)+8(x+6) = (x+8)(x+6) so (x^2+14x+48) / (x+8) = (x+8)(x+6) / (x+8) = (x+6). since x+8 cancels out in the numerator and denominator. hence x+6 is the quotient